A Gaseous compound of nitrogen and hydrogen contains 12.5% ( by mass) of hydrogen. The density of the compound relative to hydrogen is 16. The molecular formula of the compound is :

1. NH2
2. N3H
3. NH3
4. N2H4

Grade:12th Pass

1 Answers

Sunil Kumar FP
9 years ago
since a gaseous compound contain 12.5% therefore the formula is N2H4
percentage of hydrogen in NH2 is 2/16*100=12.5%
in other compund
nh3=3/17*100=17.6%
N3H=1/44*100=2.27%
N2H4=4/32*100=12.5%
DENSITY =16
therfore molecular formula is 32